OpenStudy (jessicawade):

1 question ONLY population proportion help please.

OpenStudy (jessicawade):

Use a sample n = 840, p^ = 0.25, and a 90% confidence level to construct a confidence interval estimate of the population proportion, p. A. 0.221 < p < 0.279 B. 0.214 < p < 0.286 C. 0.237 < p < 0.263 D. 0.225 < p < 0.275

OpenStudy (jessicawade):

i dont know how to find the answer and all i need is to figure this problem out

jimthompson5910 (jim_thompson5910):

one sec while I think it over

OpenStudy (jessicawade):

ive spent 2 hours on this lesson and this is the only one i cant figure out lol

jimthompson5910 (jim_thompson5910):

the formula is \[\Large L = \hat{p} - Z*\sqrt{\frac{p(1-p)}{n}}\] \[\Large U = \hat{p} + Z*\sqrt{\frac{p(1-p)}{n}}\] where L is the lower bound of the confidence interval U is the upper bound

jimthompson5910 (jim_thompson5910):

sorry typo \[\Large L = \hat{p} - Z*\sqrt{\frac{\hat{p}(1-\hat{p})}{n}}\] \[\Large U = \hat{p} + Z*\sqrt{\frac{\hat{p}(1-\hat{p})}{n}}\] ok fixed now

OpenStudy (jessicawade):

hmm lets see a=0.10 because 90% so a/2=0.05 and 1- 0.05 is 0.95 and so that on the z-chart is 1.645

jimthompson5910 (jim_thompson5910):

In this case, \[\Large \hat{p} = 0.25\] \[\Large Z = 1.645\] \[\Large n = 840\] the value Z = 1.645 is found using the normal distribution. The idea is to find the value of k such that P( -k < Z < k) = 0.90. That value of k is roughly 1.645

jimthompson5910 (jim_thompson5910):

yeah you beat me to it lol

OpenStudy (jessicawade):

lol

OpenStudy (jessicawade):

so then i plug them into the formula?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (jessicawade):

ok 1 sec

OpenStudy (jessicawade):

confused lol

jimthompson5910 (jim_thompson5910):

where at?

OpenStudy (jessicawade):

wouldnt it be E=Za/2 sqrt P^Q^/n

OpenStudy (jessicawade):

p-hat and q-hat

OpenStudy (jessicawade):

idk ive never seen the formula you used lol

jimthompson5910 (jim_thompson5910):

well q-hat is simply 1 - (p-hat)

jimthompson5910 (jim_thompson5910):

\[\Large \hat{q} = 1 - \hat{p}\]

jimthompson5910 (jim_thompson5910):

so you can think of \[\Large L = \hat{p} - Z*\sqrt{\frac{\hat{p}(1-\hat{p})}{n}}\] as \[\Large L = \hat{p} - Z*\sqrt{\frac{\hat{p} * \hat{q} }{n}}\] same for the formula for the upper bound U

jimthompson5910 (jim_thompson5910):

also, the whole portion from Z to the end of the expression is the margin of error that's why we have plus or minus that whole mess (plus or minus the margin of error)

jimthompson5910 (jim_thompson5910):

So they just said "Let E = Za/2 sqrt(P^Q^/n)"

OpenStudy (jessicawade):

ohhhhhh

jimthompson5910 (jim_thompson5910):

and I'm simplifying things. Instead of saying Za/2, I'm just saying Z

OpenStudy (jessicawade):

i got a negative lol

jimthompson5910 (jim_thompson5910):

hmm

jimthompson5910 (jim_thompson5910):

what did you get?

OpenStudy (jessicawade):

-0.021

OpenStudy (jessicawade):

xD

jimthompson5910 (jim_thompson5910):

for which part?

OpenStudy (jessicawade):

well i did 0.25-1.6545 for the 1st part of the L

OpenStudy (jessicawade):

because of P^-Z

jimthompson5910 (jim_thompson5910):

what did you get when you computed the square root of p*q/n

OpenStudy (jessicawade):

0.0149404

OpenStudy (jessicawade):

because \[\sqrt{0.25}(0.75)/840\]

jimthompson5910 (jim_thompson5910):

so you'll have 0.25 - 1.645*0.0149404 for the lower bound L

jimthompson5910 (jim_thompson5910):

I think you just calculated 0.25 - 1.645 ?

OpenStudy (jessicawade):

yeah first

OpenStudy (jessicawade):

i got 0.225423042 just now for what you said

OpenStudy (jessicawade):

so it would be 0.226? rounded

jimthompson5910 (jim_thompson5910):

0.225

OpenStudy (jessicawade):

ohhhhh

OpenStudy (jessicawade):

so that means its D?

jimthompson5910 (jim_thompson5910):

to get U, you just change the - to a + U = 0.25 + 1.645*0.0149404

OpenStudy (jessicawade):

ok

OpenStudy (jessicawade):

0.274

OpenStudy (jessicawade):

0.275*

jimthompson5910 (jim_thompson5910):

0.274576958 rounds to 0.275 so yes, D

OpenStudy (jessicawade):

yay!!

OpenStudy (jessicawade):

THANKS!! :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (jessicawade):

its confusing lol

jimthompson5910 (jim_thompson5910):

yeah confidence intervals can be tricky and odd

OpenStudy (jessicawade):

thats what i was thinking earlier lol thanks again